OpenStudy (anonymous):

Complete the square: x^2+16x+68. Please show steps...

OpenStudy (chaise):

x^2+16x+68 Half the b(64) number, square it. x^2+16x+64+4 (x+8)^2+4

OpenStudy (anonymous):

Why is b 64?

OpenStudy (anonymous):

Ohhh i need to take half and square that, got it.

OpenStudy (chaise):

If you half b, which is 16, and square it, you get 64. This is how you get a square complete. So it can be written as (a+b)^2

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

x^2+16x+64= -68+64 ===> (x+8)^2 = -4 ===> no solution in R but there two solution in complex numers , that is x+8 = 2i ==> x=-8-2i or x+8 = -2i ==> x=-8-2i